What is the mass concentration of salt ions in a solution that is made by dissolving 15 g of potassium bromide in water ?
1
Expert's answer
2017-11-27T11:21:07-0500
KBr is a compound that dissociates into cations and anions, the minimum concentration of each of those two products will be equal to the concentration of the original compound. Here's how that works:
KBr(aq)→K+(aq)+Br−(aq)
Potassium bromide dissociates into K+ cations and Br− anions when dissolved in water and 1 mole of KBr will produce 1 mole of K+ and 1 mole of Br−.
Let's assume that we have 1000 g of a KBr solution.
Mr (KBr) = 119 g/mol, so n (KBr) = 15/119= 0.126 mol
n (KBr) = n (K+)=n (Br-)=0.126 mol
m (K+)= 0.126*39 = 4.92 g
m (Br-)=0.126*80=10.08 g
Then concentration by mass of the K+ ion will be
%m (K+) =(4.92 g/1000 g)⋅100=0.492 %
%m (Br-) =(10.08 g/1000 g)⋅100=1.08 %

Answer: 0.492 % of K+, 1.08 % of Br-